# From Empirical Observations to Universality: Dynamics of Deep Learning with Inputs Built on Gaussian mixture

### v1

@@ III Method
In this section, we detail our approach to configuring Gaussian mixture inputs and explain the methodology developed to analyze the shift in dynamics as the distribution changes from a simple Gaussian to a Gaussian mixture. Our study’s dynamic metrics of interest are $Q$, $R$, and $v$.
@@ III.1 Gaussian mixture setting
In this study, instead of the simple Gaussian input ($\mathcal{N}(0,I)$) used in previous research, we employ a more generalized Gaussian mixture distribution as the input of teacher model $C$. Here, we describe the Gaussian mixture setting utilized in our analysis. Consider a random variable $r$ emerging from a $D$-dimensional Gaussian mixture with $m$ mixture components. $r$ can be represented as:
$r=r_{i}\sim\mathcal{N}(\mu_{i},\sigma_{i}^{2})\in\mathbb{R}^{D}\text{ with probability }p_{i}$
where the sum of the probabilities $\sum^{m}p_{i}=1$.
To gauge the divergence of our Gaussian mixture from a standard Gaussian distribution, we utilize the Kullback–Leibler (KL) divergence as the metric for distribution distance. The KL divergence quantifies the statistical discrepancy between two distributions, defined for the divergence of $P$ from $Q$ as:
@@
For our Gaussian mixture’s components, we fix the deviation $\sigma_{i}=1$ for all and assign the means $\mu_{i}$ by uniformly distributing them within the interval $[-\alpha,\alpha)$. Hence, a random variable $r$ from Gaussian mixture distribution comprising $m$ Gaussian components is formalized as follows:
$r=r_{i}\sim\mathcal{N}(\mu_{i},I),\quad\mu_{i}\sim\mathcal{U}[-\alpha,\alpha),\text{ with }p_{i}$ (17)
For convenience, we denote this specific Gaussian mixture distribution as $\mathcal{P}$. By defining the Gaussian mixture model in this manner, we ascertain that increasing $\alpha$ leads to a monotonically increasing KL divergence from a single Gaussian (Appendix D).
This methodology allows us to undertake empirical investigations across a spectrum of Gaussian mixtures by adjusting the parameters $\alpha$ and $m$.
III.2 Analyzing Discrepancies in Dynamics: Gaussian to Gaussian Mixture Inputs
This section explains the methods to explore how neural network dynamics deviates as input ($C$) originally assumed Gaussian, transitions towards a Gaussian mixture. Our examination centers around metrics such as $Q$, $R$, and the second-layer weight $v$ of the student model. While the generalization error is undoubtedly a critical feature, the behavior of error decreasing is a common characteristic under trainable datasets, distinguishing it slightly from $Q$, $R$, and $v$.
The methodology initiates with the assignment of initial values to the weights ($\tilde{W},\tilde{v},W,v$) and the feature matrix $F$. Firstly, with $C$ drawn from a Gaussian mixture distribution $\mathcal{P}$, we obtain $\hat{Q}_{k,\ell}(t)$, $\hat{R}_{k,m}(t)$ and $\hat{v}_{k}(t)$ at each step of the training process via SGD.
Subsequently, using identical initial values ($\tilde{W},\tilde{v},W,v$) and the feature matrix $F$, we obtain $Q_{k,\ell}(t)$, $R_{k,m}(t)$, and $v_{k}(t)$ through ODEs based on theoretical analysis (Appendix B). This is theoretically equivalent to the dynamics observed with SGD when $C$ originates from a simple Gaussian distribution $\mathcal{N}(0,I)$.
Theoretical ODE results ($Q_{k,\ell}(t)$, $R_{k,m}(t)$, $v_{k}(t)$) and empirical SGD results ($\hat{Q}_{k,\ell}(t)$, $\hat{R}_{k,m}(t)$, $\hat{v}_{k}(t)$) are collected at each time step. Given that both sets of dynamics commence from identical starting conditions, significant discrepancies are not initially observed. To evaluate the extent of divergence in network dynamics, we employ the temporal Log-sampled Root Mean Squared Error (tLog-RMSE, $\mathcal{M}$) as our metric. We temporally log-sampled both ODE results and SGD results then computed the Root Mean Squared Error (RMSE) for these sampled data points. For instance, the tLog-RMSE $\mathcal{M}$ for $v_{k}(t)$ and $\hat{v}_{k}(t)$ can be formulated as follows:
$\mathcal{M}(v_{k}(t))=\frac{1}{T}\sum_{\mathbf{t=1}}^{T}\sqrt{(v_{k}^{2}(\mathbf{t})-\hat{v}_{k}^{2}(\mathbf{t}))}$ (18)
where $\mathbf{t}$ represents the log-sampled time points derived from the original $t$ time steps. Furthermore, we examine two distinct scenarios concerning the teacher model’s input $C$: standardization and un-standardization.
In the unstandardized scenario, $C$ is directly utilized from the distribution $C\sim\mathcal{P}$. Conversely, the standardized scenario implements a simple standardization method, where the random variable $C$ is rescaled as
$C:=(C-\mathbb{E}[C])/\sqrt{\mathbb{E}[(C-\mathbb{E}[C])^{2}]}.$ (19)
The expectation value is computed by empirical samples, $\{C\}$. For the sake of clarity and convenience in notation, we refer to the standardized setting as $C\sim\overline{\mathcal{P}}$.
III.3 Additional Detailed Experimental Conditions
In this study, the $C$ dimension was set to $D=500$ and the $X$ dimension for student input to $N=1000$. The dimensions of the hidden layers for both teacher and student models were uniformly set to $K=M=2$. For the activation functions, both the teacher and student models utilized the same function, $g(x)=\tilde{g}(x)=\text{erf}(x/\sqrt{2})$ or ReLU function. The nonlinear function $f(x)=\text{sgn}(x)$ was employed to generate the student input. The learning rate was set at $\gamma=0.2$, and training was conducted using the quadratic loss $\mathcal{L}=1/2(\hat{y}-y)^{2}$. The SGD update used was a single batch SGD with layerwise learning rate scaling, as mentioned earlier II.2. For training, the neural network was updated for a total of $100\times 1000$ steps.
IV Results
In this section, we delve into the empirical exploration of how the dynamics shift as the input distribution $C$ transitions from a simple Gaussian to a Gaussian mixture. Through our investigation, we evaluated 20 distinct Gaussian mixtures, each undergoing $100\times 1000$ training steps. Since SGD dynamics are influenced by randomness and ODE dynamics can fluctuate by numerical errors, it is possible for the results of SGD in the simple Gaussian setting to differ from those of the ODE results, that should theoretically be identical. Out of the 20 different Gaussian mixtures, excluding those where the ODE results and the SGD dynamics of the simple Gaussian significantly different, we sampled 10 for analysis.
The results provided the average tLog-RMSE value across these varied mixtures. The dimensions of matrices and vectors are $Q,R\in\mathbb{R}^{2\times 2}$, and $v\in\mathbb{R}^{2}$.
For instance, since the dynamic metric $Q$ consists of a total of four values ($Q_{k,\ell}$), the average tLog-RMSE value $\mathbb{E}[\mathcal{M}(Q_{k,\ell})]$ is calculated as follows:
$\mathbb{E}[\mathcal{M}(Q_{k,\ell})]\equiv\mathcal{M}(Q)=\mathbb{E}_{\{k,\ell\}}[\mathcal{M}(Q_{k,\ell}(t))]$ (20)
As defined above, for convenience of notation, the absence of an index indicates the average, while the explicit presence of an index refers to the tLog-RMSE value for that specific index.
We calculate the KL divergence $D_{\text{KL}}(\mathcal{P}\|\mathcal{N})$, across different $\mathcal{P}$ that parametrized by $\alpha$ as mixture setting outlined earlier (17). It then assessed the influence of these $D_{\text{KL}}(\mathcal{P}|\mathcal{N})$ values on the discrepancies, tLog-RMSE. Note the spectrum of KL divergence varies across different components, depending on $\alpha$. In our Gaussian mixture setting, since the components have mean values $\mu\in[-\alpha,\alpha)$, an increase in the number of components can lead to overlapping phenomena, thereby reducing the KL divergence. This effect can be observed in subsequent results, where slightly different lengths along the KL divergence axis are shown.
In the main text, our emphasis is on the results obtained using the ReLU activation function. For outcomes related to the erf activation function, refer to Appendix G.
IV.1 Results from un-standardized Gaussian mixture
Figure 2: Example of Dynamics under un-standardized Gaussian mixture with $m=2$, and $\alpha=9$. From left to right, the plots show the dynamics of generalization error $\epsilon_{g}$, covariances $Q$ and $R$, and the weight of second layer $v$. The shaded area represents the saturation region, indicating the stage where training is nearing completion, with the error remaining small and unchanged. The SGD results are averaged over 10 runs.
As observed in Figure 2, the dynamics under un-standardized Gaussian mixture settings significantly diverge from those under a simple Gaussian distribution. Analyzing various mixture settings, we investigate the relationship between KL divergence and the error metric $\mathcal{M}$.
Figure 3: Log Sampled MSE under KL divergence in un-standardized Gaussian mixture setting, under ReLU activation function. a, b, and c display the results corresponding to metrics $Q$, $R$, and $v$, respectively. The shaded areas around each curve indicate the standard error, across varying numbers of Gaussian mixture components $m$.
Figure 3 intriguingly demonstrates that despite significant deviation from $\mathcal{N}(0,I)$ (increasing KL divergence), certain dynamic metrics— $Q$ and $v$—tend to converge towards an asymptotic limit. This results simply explained by non-divergence (or non-explosive) nature of our learning dynamics. Despite the Gaussian mixture deviating significantly from a simple Gaussian distribution, the student’s trainable parameters do not explode. As a result, the non-divergence property guarantees that the error in the dynamics metric does not exhibit exponential behavior.
IV.2 Results from standardized Gaussian mixture
Subsequently, we investigated the dynamics under teacher model inputs $C\sim\overline{\mathcal{P}}$, derived from the standardization of Gaussian mixture samples.
Figure 4: Example of Dynamics under standardized Gaussian mixture with $m=2$, and $\alpha=9$. From left to right, the plots show the dynamics of generalization error $\epsilon_{g}$, covariances $Q$ and $R$, and the weight of second layer $v$. Except for the saturation region, where the effects of randomness are high, the dynamics almost align with conventional predictions (ODE or SGD under Gaussian). The SGD results are averaged over 10 runs.
Interestingly, the dynamics under standardized Gaussian mixture settings closely resemble those under a simple Gaussian distribution. In Figure 4, this similarity is readily apparent.
Figure 5: Temporal log sampled RMSE(tLog-RMSE) under KL divergence in standardized Gaussian mixture setting, under ReLU activation function. a, b, and c display the results corresponding to metrics $Q$, $R$, and $v$, respectively, with the standard error represented by the shaded regions around each curve, across different numbers of Gaussian mixture components $m$. To illustrate the minimal scale of the error, a dotted line representing the mean of the each dynamic metric values $\mathbb{E}[Q]$, $\mathbb{E}[R]$, $\mathbb{E}[v]$ is added, highlighting the small error magnitude in comparison to the overall scale of the values. Additionally, a randomly sampled example of evolution of dynamic metrics, demonstrating almost negligible error, is included as an inset.
The analysis of dynamics with standardized Gaussian mixtures yielded notably insightful outcomes. As illustrated in Figure 5, the average tLog-RMSE $\mathcal{M}$ for various mixtures appears minimal relative to the metrics’ scale, indicated by a dotted line. These mixtures do not introduce any significant discrepancies between theoretical predictions and empirical observations. The subsequent Discussion section V will provide explanation of these phenomena, why mixture based experiment result conceptually converge with conventional theory.
@@ V Discussion
In this section, we delve into the mathematical analysis of the convergence properties of the standardized Gaussian mixture. As discussed in the Background section II, the dynamics of our teacher-student model setting are influenced by the distribution followed by $\{\lambda,\nu\}$. Therefore, to analyze how network dynamics change when $C$ is a Gaussian mixture, it is crucial to examine the distribution of $\{\lambda,\nu\}$ under such a mixture setting.
Following the proof sequence for the Gaussian Equivalence Property II.3 where ${\lambda,\nu}$ adhere to a Gaussian distribution, we first investigate how the expectation of function correlations converges in the context of a Gaussian mixture, then derive a modified version of GEP applicable to this scenario.
@@ V.1.1 Correlations between Functions with Gaussian mixture
Consider $I+J$ random variables represented as $x=(x_{1},x_{2},\cdots,x_{I})^{\top}$ and $y=(y_{1},y_{2},\cdots,y_{J})^{\top}$. Each variable $x_{i}$ and $y_{j}$ originates from a Gaussian mixture: $x_{i}=X_{k}$ with probability $p_{k}$ where $X_{k}\sim\mathcal{N}(\mu_{k},\sigma_{k}^{2})$, and similarly, $y_{j}=Y_{l}$ with probability $p_{l}$ where $Y_{l}\sim\mathcal{N}(\mu_{l},\sigma_{l}^{2})$.
Despite utilizing Gaussian mixtures, each random variable can essentially be considered as following a single Gaussian distribution with a certain probability, allowing us to straightforwardly implement the existing function correlation approximation II.1. In the Gaussian mixture setting, we derive the following lemma for function correlation approximation under mixture distribution:
Lemma V.1 (Function correlation approximation with mixture distribution).
For the $I=J=1$ case with two Gaussian mixture variables $u_{1}$ and $u_{2}$ standardized to mean zero and variance one, and assuming weakly correlated covariance ($\mathbb{E}[u_{1}^{2}]=1$, $\mathbb{E}[u_{2}^{2}]=1$, $\mathbb{E}[u_{1}u_{2}]=\epsilon m_{12}$), the approximation of function correlation for the Gaussian mixture in the limit as $\epsilon\to 0$ is given by:
$\mathbb{E}[f(u_{1})g(u_{2})]=\sum_{\{p_{k},p_{l}\}}p_{k}p_{l}[\braket{f(u_{1})}_{k}\braket{g(u_{2})}_{l}]+\mathcal{O}(\epsilon)$ (21)
@@
$\braket{f(u_{1})}_{k}=\mathbb{E}_{u_{1}\sim\mathcal{N}_{k}}[f(u_{1})],\quad\braket{g(u_{2})}_{l}=\mathbb{E}_{u_{2}\sim\mathcal{N}_{l}}[g(u_{2})]$
Therefore, even with Gaussian mixtures, an approximation of function correlation can achieve a similar form.
V.1.2 Dominance of Moments in the Expectation Value of Functions
In deriving the original Gaussian Equivalence Property II.3 we utilized the approximation form of function correlation II.1 to determine the covariance matrix of $\{\lambda,\nu\}$. Since the Lemma V.1 has equivalent form with Lemma II.1, remaining question pertains to how the expectation values of functions ($\mathbb{E}[f(u_{1})]$ or $\mathbb{E}[g(u_{2})]$) under random variables $\{u_{1},u_{2}\}$, following a Gaussian mixture distribution, differ from those assuming random variables $\{u_{1},u_{2}\}$ following simple Gaussian distribution.
In our study, we particularly employed the sgn function as our function $f$ within the $\lambda=f(U)W^{\top}/\sqrt{N}$. Dissecting the sgn function into differentiable regions reveals that higher-order derivative terms become negligible. This insight, coupled with Taylor expansion, allows us to probe the characteristics of the expectation value $\mathbb{E}[f(x)]$ for a random variable $x$ following an arbitrary distribution.
For simplicity, we introduce the following notation:

### v2

@@ III Method
In this section, we detail our approach to configuring Gaussian mixture inputs and our experimental settings, where we consider a two-layer teacher-student model as in Eq. 2.
@@ III.1 Gaussian Mixture Setting
Figure 2: Kullback–Leibler (KL) divergence $D_{\text{KL}}(\mathcal{P}||\mathcal{N})$ for different values of $\alpha$ and number of components $m$.
In this study we employ a generalized Gaussian mixture distribution as the teacher model input, $C$. Below, we describe the specific Gaussian mixture setting utilized in our analysis. For each Gaussian mixture component, we fix the covariance matrix $\Sigma_{i}=I$, where $I$ is the identity matrix, and assign the means $\mu_{i}$ by uniformly distributing them within the interval $[-\alpha,\alpha)$. The weights $p_{i}$ are randomly selected from a uniform distribution and normalized so that $\sum p_{i}=1$. A random variable $r$ from the Gaussian mixture distribution, comprising $m$ Gaussian components, is formalized as follows:
$r=r_{i}\sim\mathcal{N}(\mu_{i},I),\quad\mu_{i}\sim\mathcal{U}[-\alpha,\alpha),\text{ with }p_{i}.$ (17)
For convenience, we denote this specific Gaussian mixture distribution as $\mathcal{P}$. This methodology allows us to conduct empirical investigations across a spectrum of Gaussian mixtures by adjusting the parameters $\alpha$ and $m$. In Fig. 2, we check that increasing $\alpha$ leads to a monotonically increasing Kullback–Leibler (KL) divergence from a single Gaussian distribution. We estimate KL divergence by using $k$-nearest-neighbor distances [37].
III.2 Additional Experimental Settings
The dimension of teacher model input $C$ was set to $D=500$, and the dimension of student model input $X$ was set to $N=1000$. The dimensions of the hidden layers for both teacher and student models were uniformly set to $K=M=2$. Both the teacher and student models employed the same activation function, $g(x)=\widetilde{g}(x)=\text{erf}(x/\sqrt{2})$ or ReLU. The nonlinear function $f(x)$ used to generate the student input was either $f(\cdot)=\text{sgn}(\cdot)$ or $f(\cdot)=tanh(\cdot)$.
The learning rate was set to $\gamma=0.2$, and training was conducted using the quadratic loss $\mathcal{L}=\frac{1}{2}(\hat{y}-y)^{2}$ with a scaled SGD update procedure as shown earlier (Eq. (4)). The student model was updated for a total of $100\times 1000$ steps. We used the same initial conditions ($\widetilde{W},\widetilde{v},W,v$) and the feature matrix $F$ setting to obtain the evolution of the dynamics.
Figure 3: Examples of dynamics under unstandardized Gaussian mixtures with $m=2,m=16$, and $\alpha=0.001,0.01,0.1,1,10$. Dynamics of (a) generalization error $\epsilon_{g}$, (b) covariance matrix $Q$, (c) covariance matrix $R$, and (d) weight of the second layer $v$. The SGD results are averaged over 10 runs.
Figure 4: Difference in $v$ for various $\alpha$ values. $v_{k,\text{ODE}}$ refers to $v_{k}$ values from ODE dynamics, and $v_{k,\text{GM}}$ refers to $v_{k}$ values from SGD dynamics under the Gaussian mixture setting. The error due to simple randomness was corrected. Here, $\langle\cdot\rangle_{k}$ denotes the average over $k=1,2$, and $t$ denotes the training steps for simplicity.
Figure 5: Examples of dynamics under standardized Gaussian mixtures with $m=2,m=16$, and $\alpha=0.001,0.01,0.1,1,10$. Dynamics of (a) generalization error $\epsilon_{g}$, (b) covariance matrix $Q$, (c) covariance matrix $R$, and (d) weight of the second layer $v$. The SGD results are averaged over 10 runs.
Figure 6: KL divergence $D(\overline{\mathcal{P}}||\mathcal{N})$ for different $\alpha$ values and number of components $m$. Note that $\overline{\mathcal{P}}$ denotes the standardization setting.
IV Results
In this section, we explore how the dynamics evolve as the input distribution $C$ transitions from simple Gaussian to a Gaussian mixture. In Section II we established that when $C$ follows a simple Gaussian distribution, we can analytically trace the dynamics using the order parameters $Q$, $R$, and $T$ alongside ODEs. The results derived from these ODEs are analytically consistent with those obtained through SGD under a simple Gaussian $C$.
Here, we compare the dynamics derived from ODE computations with those from SGD across various scenarios where $C\sim\mathcal{P}$. For simplicity, we use “ODE dynamics” to refer to the ODE simulation results, which are essentially equivalent to SGD under a simple Gaussian $C\sim\mathcal{N}$, and use “SGD dynamics” to denote the results obtained from running SGD with given $C\sim\mathcal{P}$. Our primary focus is on the results obtained using the ReLU activation function and the sgn nonlinear function. For outcomes related to the erf activation function or the $tanh$ nonlinear function, refer to Appendix B.
@@ IV Results
IV.1 Unstandardized Gaussian Mixture Results
First, we examine how the dynamics deviate when the inputs have various Gaussian mixture properties. We consider $\alpha=0.001,0.01,0.1,1,10$ and $m=2,16$ for visualization. As observed in Fig. 3, the dynamics under unstandardized Gaussian mixture settings significantly diverge from those under a simple Gaussian distribution. Although both the ODE dynamics and SGD dynamics start from identical initial conditions, significant discrepancies emerge over time. As $\alpha$ increases, the divergence becomes more pronounced, matching the KL divergence trends shown in Fig. 2, where $\alpha$ exhibits greater influence on KL divergence than does the component number $m$.
Even for an arbitrary distribution, the dynamics of the second layer’s weight $v$ (Eq. (5)) maintain the same form but differ in the calculation of the function correlation. By tracking the difference in $v$ between ODE and SGD dynamics, we can assess the effect of distribution differences.
In Fig. 4, we can see how the mean difference of $v$ deviates over time; as expected, the deviation becomes more distinct as time evolves. The deviation is small when $\alpha$ is sufficiently small because the means $\mu_{i}$ for the Gaussian mixture are mostly sampled at 00, making it difficult to distinguish between samples from the Gaussian mixture and simple Gaussian distribution. This convergence due to distribution similarity occurs in the $\alpha\leq 0.1$ regime, where the KL divergence is on the order of $\mathcal{O}(10^{-2})$ (Fig. 2). Conversely, larger divergence in the dynamics is observed for large $\alpha$, as increasing $\alpha$ enhances the dissimilarity between the Gaussian mixture and simple Gaussian distribution.
IV.2 Standardized Gaussian Mixture Results
Second, we investigated the dynamics when the teacher model inputs were standardized, i.e., $C\sim\overline{\mathcal{P}}$. As before, we consider $\alpha=0.001,0.01,0.1,1,10$ and $m=2,16$ for visualization. Analysis of the dynamics with standardized Gaussian mixtures yielded notable outcomes. As shown in Fig. 5, these mixtures do not introduce any significant discrepancies between ODE dynamics and SGD dynamics.
Further investigation into the KL divergence of $\overline{\mathcal{P}}$ reveals that this convergence phenomenon is not solely due to the similarity of the distributions. From the previous unstandardized results, we observed that the dynamics converge when the KL divergence is on the order of $10^{-2}$. However, as shown by the standardized results in Fig. 6, the KL divergence reaches an order of $10^{-1}$ in large $\alpha$ scenarios, yet the dynamics under these large $\alpha$ values still exhibit strong convergence. Even for very large $\alpha$ or a regime with large $m$, the convergence phenomena are maintained.
These results suggest that standardization plays a more crucial role in the observed convergence phenomenon beyond simply making the distribution Gaussian-like. Section V provides an explanation of this phenomenon and why the mixture-based experimental results converge with conventional theory.
@@ V Discussion
In this section, we mathematically analyze the convergence properties of the standardized Gaussian mixture. As discussed in Section II, the dynamics of our teacher-student model are influenced by the $\{\lambda,\nu\}$ distribution. Therefore, to analyze how network dynamics change when $C$ is a Gaussian mixture, it is crucial to examine the distribution of $\{\lambda,\nu\}$ under such a mixture setting.
Following the proof sequence for the GEP (Property II.3), where $\{\lambda,\nu\}$ adhere to a joint Gaussian distribution, we first investigate how the function correlation is approximated in the context of a Gaussian mixture, and then derive a modified version of the GEP applicable to this scenario.
@@ V.1.1 Correlations between functions with Gaussian mixture
Consider $I+J$ random variables represented as $x=(x_{1},x_{2},\cdots,x_{I})^{\top}$ and $y=(y_{1},y_{2},\cdots,y_{J})^{\top}$. Each variable $x_{i}$ and $y_{j}$ originates from a Gaussian mixture, $x_{i}=X_{k}$ with probability $p_{k}$ where $X_{k}\sim\mathcal{N}(\mu_{k},\sigma_{k}^{2})$, and similarly, $y_{j}=Y_{l}$ with probability $p_{l}$ where $Y_{l}\sim\mathcal{N}(\mu_{l},\sigma_{l}^{2})$.
Despite utilizing Gaussian mixtures, each random variable can essentially be considered to follow a single Gaussian distribution with a certain probability, allowing us to straightforwardly implement the existing function correlation approximation Lemma II.1. In the Gaussian mixture setting, we derive the following lemma for function correlation approximation.
Lemma V.1 (Function correlation approximation with Gaussian mixture).
For the $I=J=1$ case with two Gaussian mixture variables $u_{1}$ and $u_{2}$ standardized to mean 00 and variance $1$, and assuming weakly correlated covariance ($\mathbb{E}[u_{1}^{2}]=1$, $\mathbb{E}[u_{2}^{2}]=1$, $\mathbb{E}[u_{1}u_{2}]=\epsilon m_{12}$), approximation of the function correlation for the Gaussian mixture in the limit $\epsilon\to 0$ is given by:
$\mathbb{E}[f(u_{1})g(u_{2})]=\sum_{\{p_{k},p_{l}\}}p_{k}p_{l}\braket{f(u_{1})}_{k}\braket{g(u_{2})}_{l}+\mathcal{O}(\epsilon)$ (18)
@@
$\braket{f(u_{1})}_{k}=\mathbb{E}_{u_{1}\sim\mathcal{N}_{k}}[f(u_{1})],\quad\braket{g(u_{2})}_{l}=\mathbb{E}_{u_{2}\sim\mathcal{N}_{l}}[g(u_{2})].$
Therefore, even with Gaussian mixtures, an approximation of the function correlation can achieve a similar form as Lemma II.1.
V.1.2 Dominance of cumulants in the expectation values of functions
In deriving the GEP (Property II.3), the approximated form of the function correlation (Lemma II.1) is utilized to determine the covariance matrix of $\{\lambda,\nu\}$.
Since Lemma V.1 has a form equivalent to Lemma II.1, the remaining question pertains to how the expectation values of the functions (e.g., $\mathbb{E}[f(u_{1})]$, $\mathbb{E}[g(u_{2})]$) under random variables ($\{u_{1},u_{2}\}$) following a Gaussian mixture distribution differ from those assuming random variables following a simple Gaussian distribution.
In our study, we employed sgn in particular as our function $f$ within $\lambda=f(U)W^{\top}/\sqrt{N}$. Dissecting the sgn function into differentiable regions reveals that higher-order derivative terms become negligible. This insight, coupled with Taylor expansion, allows us to probe the characteristics of the expectation value $\mathbb{E}[f(x)]$ for a random variable $x$ following an arbitrary distribution.
For simplicity, we define the following notation.